You are an analyst employed by a yacht manufacturer that last year sold 30,000 luxury yachts at $500,000 each. Your market research indicates that:


i) the price elasticity of demand for your aircrafts in −0.5. (or +0.5 in absolute value); 

ii) the income elasticity of demand for your aircrafts is +2.6; and 

iii) the cross price elasticity for your aircrafts with respect to the price of a comparable jet manufactured by a competitor is +1.4


Answer the following questions based on the information above:


A. Suppose that you expect a ceteris paribus decrease in average incomes of 15% this year compared to last year. How many aircrafts do you estimate that your company will sell this year? How will it impact total revenues? (6 MARKS)


B. Assume now that you do not think incomes will change, but that you expect your competitor will decrease his price by 3%. Assuming that your company does not change the price of its aircrafts, how many would you expect your company will sell this year? (5 MARKS)



1
Expert's answer
2022-04-04T14:41:27-0400

a) The quantity sold will decrease by 15×2.6=39%, or by 30,000×0.39=11,700 yachts till 30,000×(1-0.39)=18,300 yachts. This will lead to the revenue decrease by 39%, or 11,700×500,000=5,850 million dollars

b) The quantity sold will decrease by 3×1.4=4.2%, or 30,000×0.042=1,260. So, the company will sell 30,000-1,260=28,740 yachts